KAMI Shift Summary for Owl 1-30-00


Subject: KAMI Shift Summary for Owl 1-30-00
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:44:59 CST


           KAMI-150 Shift Summary for Owl shift, 30/Jan/00

On Shift: Y.B. Hsiung (with Wah, RickK, AnatoliL early on)

Hours of beam delivered at all intensities: 5.8 out of 7.0
Hours of beam delivered at 2E11 intensity: 5.0
Hours of any KAMI test data: 5.8

Accounting of lost time: TeV quenches.

========================================================================
Runs ended this shift:
Run 15684 - KAMI_Hadron (no disk, no tape)
            1st 150 GeV KAMI hadron beam test run. ~230 KL->3pi0.
Run 15685 - KAMI_Muon (long muon run in progress) ~838K at end of shift.
            low intensity 2E11 and very low muon rate.

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
   KAMI 150 GeV beam on target with 20 mrad targeting angle and 20" Be
absorber. We took ~40 spills 1st 150 GeV hadron beam to check out
UCAL, CsI and L3 etc. Got 1st UCAL beam hadron energy spectrum (see
KAMI_LOG_1 pp170) and KL->3pi0 (pp172-173) from online plots.
Switch back to muon and taking a long muon run. Four TeVatron quenches
during the shift since the 150 GeV beam tune and extraction were still
not ideal. (Beam study will continue during the day shift, hopefully
will be much better later)

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
              00:00 Had TeV quench, wait for recovery.
15684 00:10 start KAMI_Hadron (no disk, no tape) as a test
                       call MCR to open the beam stop, NM2S1=275A, all
                       other magnets = 0A.
              00:20 1st KAMI hadron beam dribbled in, hardly noticed
                       by the audio spill monitor, except online histos.
                       Intensity ~2.5E11 to us, but 8E11 in TeV, large
                       loss somewhere.
              01:44 Another quench, est. 45 min.
15684 02:00 Stop run, close beam stop for muon.
15685 02:11 beam return for long muon run.
                       change the audio spill monitor prescaler from
                       PS=999 to 20 so we can hear the beam spill.
              04:57 Quench at C4, est. 45 min.
              05:08 Beam return to KAMI.
                       Chat with MCR about the reason of quench and
                       they said there are beam instabilities because
                       the tune and QXR is not ideal yet. They will
                       work on this during today's day shift.
              06:50 Another quench at C4, est. 45 min.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Ignore red SDAQ buttons "OnSplit" and "VolMon2 (UCLA stuff)"

Before the beam study, stop the current long muon run, then
start another one.



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:45:02 CST